Pet Place Supplies Inc., a pet wholesale supplier, was organized on May 1. Projected sales for each of the first three months of operations are as follows:

May $310,000
June 420,000
July 560,000

All sales are on account. Of sales on account, 51% are expected to be collected in the month of the sale, 45% in the month following the sale, and the remainder in the second month following the sale.

Prepare a schedule indicating cash collections from sales for May, June, and July.

Pet Place Supplies Inc.
Schedule of Cash Collections from Sales
For the Three Months Ending July 31
May June July
May sales on account:
Collected in May $
Collected in June $
Collected in July $
June sales on account:
Collected in June
Collected in July
July sales on account:
Collected in July
Total cash collected $ $ $

If you could provide explanation, I would greatly appreciate it!
Answer #1
Pet Place Supplies Inc.
Schedule of Cash Collections from Sales
For the Three Months Ending July 31
May   June july
May sales on account:
Collected in May $310,000*51% = $158,100
Collected in June $310,000*45% = $139,500
Collected in July $310,000*4% = $12,400
June sales on account:
Collected in June $420,000*51% = $214,200
Collected in July $420,000*45% = $189,000
July sales on account:
Collected in July $560,000*51% = $285,600
Total cash collected $                                  1,58,100 $                                  3,53,700 $                         4,87,000

You can reach me over comment box, if you have any doubts. Please rate this answer